UT Sep 11, 2019 - Restart Night #10


OSA: Josh Williams

Observers: John, Olga

Summary

Clouds and fog, and there was no chance of opening.

Collected data for LUCI1 + LUCI2 flexure compensation lookup tables for these modes:
L1 L2 Comments
L1_N375_G150 L2_N375_G040 OK, not too many spots with Ar for N375+G150 @ Ks, voltage = -0.6 for L2
L1_N180_MIRR L2_N180_MIRR spots well detected
L1_N375_G150 L2_N30_G040 repeated L1 config, this time with Ar+Kr+Xe, L2 spots weak, may need longer exptime

Log

Clouds, some precipitation. Looks like there will be no chance to open soon.

02:50 Set up for collection of LUCI flex comp data.
  • L1: N3.75 + G150 (Ks)
  • L2: N3.75 + G040 (J and need to set voltage to get spectra on detector)

Work on L2 configuration:
  • luci2.20190911.0022 (03-22) is first halo2 spectrum using the voltage corresponding to central wavelength 1.25 mic. HIRAMO says this voltage is -1.29
  • set target voltage to -1.0 and take next set of images (save mode normal, ndit 20, so 23-42)...
  • repeat with savemode integrated.
    • 43 voltage = -1.0 (DIT x NDIT x NEXPO = 2.51 x 20 x 1) J and HALO2
  • set target voltage to -0.5 (loop remains closed)
    • 44 voltage = -0.5 J
      • looks better than 43 - the traces at the bottom of the detector shift to the right and are fully on the detector.
    • 45 voltage = -0.6 J @ 1.25 mic (-0.6 is what Dave read for the central wavelength for 2.2 mic)
    • change to K & 2.2 mic, just to see. Target voltage went to -0.57, but setting it back to -0.60.
    • 46 voltage = -0.60 K (traces are covering the left 3/4 of the detector).
    • 47 voltage = -0.40
    • 48 voltage = -0.20
    • 49 voltage = 0.0
    • 50 voltage = 0.2
    • 51 voltage = 0.4
    • 52 voltage = 0.6
    • 53 voltage = 0.8
    • 54 voltage = 1.0
    • 55 voltage = -0.6 with HKspec (vs K)
    • 56 voltage = 0.6 with HKspec (vs K)
    • 57 voltage = 0.6 with K (no, because I switched filter, voltage went back to -0.57)
    • 58 voltage = 0.6 with K (should be like 52).
  • L1 Ar spectrum with N3.75 @ Ks, central wavelength 2.15 mic.
    • L1 optspec sieve + G150 @ Ks with Argon. 2.51s x 2 too little (0002)
    • L1 3,4 uses dit x ndit = 30s x 2 (ok, see spots)
    • L1 5 halo 2 20s x 1 x 1
  • Seeing a gradient from bottom to top - Josh is homing the AGWs.
    • 59 voltage = 0.6 with K (gradient gone)
    • 60 voltage = 0.6 with J (switch to dit x ndit = 2.51s x 2)
  • with J & 1.25 mic, step through voltage with 0.5 volt steps
    • 61 voltage = 1.0
    • 62 voltage = 1.5
    • 63 voltage = 2.0
    • 64 voltage = 2.5
    • 65 voltage = 1.0 (open loop, set, close loop)
    • 66 voltage = 0.5
    • 67 voltage = 0.0
    • 68 voltage = -0.5
    • 69 voltage = -1.0
    • 70 voltage = -1.50 moving to the left and pretty blank on the right now
    • 71 voltage = -2.0
    • 72 voltage = -2.5 way off to the left now...
  • analyze results: 61 through 72 ... scale from 0-300 so see only brightest traces. Appears that 0.5 (66) is best.
  • Go to 0.6 (best before with higher res steps).
    • 73 voltage = 0.6 J 1.25mic
    • 74 voltage = 0.6 K 2.2 mic (just for fun)
    • 75 voltage = 0.6 J 1.25 mic again (but voltage was -1.29 because of change of config)
    • 76 voltage = 0.6 J 1.25 mic (this time volage set)
  • Argon
    • L1 8 and L2 77 (lots and lots of spots for L2, not so many but also spots for L1).
  • Going to start flex comp run:
    • Josh is moving the rotators to -90 deg.
    • at zenith
    • L1:
    • L2:
      • N375
      • OptSpec _Sieve
      • G040AO @ 1.25 mic (voltage = 0.60)
      • clear + J
      • FlexComp OFF
      • Calib in, AGW probe out of beam
      • Argon ON
    • Starting ...
tel L1 L2
zen 09-81 78-150
68 deg 82-154 151-223
46 deg 155-227 224-296
24 deg 228-300 297-369
5s darks 301-303 370-372

  • Now N180 + MIRR for both sides. Use thermal emission. Switching to N180, I noticed error messages:
2019.09.11 07:34:47 s | error high     | luci.luc | Luci ONE        | ONE_Compensatio | AdjustMirrorThread.java#adjustToZeroPoin(475)    | Could not calculate actual shift to zero-point position for camera=0 grating=1
2019.09.11 07:34:47 s | error lowest   | luci.luc | Luci ONE        | ONE_Compensatio | AdjustMirrorThread.java#adjustToZeroPoin(407)    | Previous camera or grating position was not registered in the flexure compensation unit. Actual will be used for alignment.

On L1, toggle from mirror to G150Ks and back. Same error. Again on going from G150Ks to mirror with N1.8. Move to N3.75 and try. Same thing when moving camera and then when moving from G150Ks to mirror with N3.75. Take a test image with optspec_sieve and miror.

L1 304 is the test image with N375 and mirror. The spots look pretty well centered. Unsure about the reason for the error message. It seems to indicate that a zeropoint shift for FM4 cannot be calculated. Is there a table in place? or will it not do this if FC is OFF? We want FC OFF and the field appears to be pretty well aligned. (1023,1034). Continue despite error message.

L2 373 is a test image with N375 and mirror. Field appears to be very well aligned. (1024, 1023).

Test images with N1.8: 305 and 374

  • L1 and L2:
    • N180
    • OptSpec _Sieve
    • Mirror
    • clear+K
    • Flex Comp OFF
    • Calib IN
    • No lamps - ambient light.

Started at 07:55. Spots look pretty aberrated on L2. Noticing voltage loop is still closed, but this should not matter since mirror is in use on L2, right?
24 deg 306-378 375-447
46 deg 379-451 448-520
68 deg 452-524 521-593
90 deg 525-597 594-666
667 one L2 image with voltage loop open - looks similar.

  • Now L2 N30+G040 clear + J Argon
    • 668 test image with default voltage, -1.29 - no lamp
    • 669 test with Argon and voltage = -1.29 - see spots but faint and perhaps not well centered, as expected.
    • 670 test with Halo2 and -1.29 volts. Nothing. Try Halo1 (loop was open, also)
    • 671 Halo1 N30+G40 5s x 1 x 1. (loop open) - see traces.
    • 672 Halo1 N30+G040 5s x 1 x 1 voltage = 0.6 loop closed.
    • 673 Argon N30+G040 5s dit voltage = 0.6
    • 674 Halo1 N30+G040 5s x 1 x 1 voltage = 0.6
    • 675 same but voltage = 1.1
    • 676 same but voltage = 0.1
    • 677 voltage 0.6 and Xe(faint) (probably 0.5 is too large a step size for N30. Going back to 0.6 as spots were seen).
    • 678 Ar and 0.6
  • On L1, try again N375+G150Ks
    • 598 One lamp only (Xe)
    • 599 Ne+Ar+Xe for more spots
  • L2:672
    • N30
    • OptSpec _Sieve
    • G040 @ 1.25
    • clear+J
    • FC OFF
    • Argon
  • L1:

zen 600-672 679-751
68 deg 673-745 752-824
46 deg 746-818 825-897
24 deg 819-891 898-970
Done. I'm not sure the N30+G040 data will be of any use. The spots are visible, but they are not well detected and iraf.imexam cannot centroid on the few I checked. Plus they seem to shift quite a lot. But the N375+G210Ks may have been worth repeating with all lamps, Ar+Kr+Xe, on.

13:28 UT. LUCI1 and LUCI2 in end-of-night config with mirror, blinds, optspec_sieve and cal unit out. Lamps all off

Outline of Plan

Plan rolled over from Night #7-8-9
  • LUCI
    • AGW-1 Transform
      • No need to update PCS, rotator ZP looks good after astrometric mask test
    • IF spectroscopic modes are released for LUCIs
      • G200 science-like OB
      • G210 science-like OB
  • Test patch to GCS/PCS to address dithering NS guiding
    • LUCI
    • MODS
  • DX M3 alignment
  • LBC
    • Standard photometric fields - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters
    • LBC sub-framing (test OBs needed)
    • NS guiding with dithers
    • Checkout observations with Bessel filters
  • MODS
    • Test NS guiding with dithering
Closed Dome tests:

  • Collection of LUCI flex comp data: Dave re-reduced all of the flex comp data with the corrected offsets and updated these to the database.
  • Modes remaining:
L1 L2 Comments
L1_N375_G150 L2_N375_G040 (third gratings w/N375)
L1_N180_MIRR L2_N180_MIRR (N180 imaging)
  L2_N30_G040 (N30_G040)
L1_N180_G150 L2_N180_G040 (third gratings w/N180)
  L2_N30_G210 (N30_G210)

-- BarryRothberg - 10 Sep 2019
Topic revision: r6 - 11 Sep 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback